OpenStudy (anonymous):

One circular table has a diameter of 8 ft, and another circular table has a diameter of 12 ft. How much greater is the area of the larger table? Round to the nearest whole number.

OpenStudy (poetryprincess):

Pi(4)^2 = 50 Pi(6)^2 = 113 113-50 = 63. The answer is 63.

OpenStudy (poetryprincess):

@yfugate